A stock is expected to pay dividends of $1.00, $0.75 and $2.00 for the next 3 years, respectively. After that dividends are expected to grow at a constant rate of 6% indefinitely. The required return on the stock is 10%. Compute the present value of the non-constant dividends.
Answer #1

Step 1: Computation of market price at the end of year 3 using Gordon Growth Model

P3 = D4 / (Ke – g)

Where,

P3 - Market price at the end of year 3 =?

D4 - Expected dividend in year 4 = 2*1.06 = 2.12

Ke – Cost of equity = 10%

G – Growth rate in dividend = 6%

P3 = 2.12/(.10-.06)

= 2.12/.04

= 53.00

Step 2: Computing current share price by discounting the cashflow at required return

Year Dividend PVF@10% Present Value (Cashflow*PVF)
1 1            0.909 0.91
2 0.75            0.826 0.62
3 2            0.751 1.50
3 53 0.751 39.82

current share price = Cashflow*PVF

= .91+.62+1.50+39.82

= 42.85

present value of the non-constant dividends = .91+.62+1.5

= 3.03

You can use the equation 1/(1+i)^n to find PVF using calculator